OpenStudy (anonymous):

evaluate the indefinite integral. (1/sqrtx)dx

OpenStudy (turingtest):

think of how to represent \[\sqrt x\]in the form\[x^a\]

OpenStudy (anonymous):

x^1/2

OpenStudy (turingtest):

so\[{1\over\sqrt x}=?\]

OpenStudy (turingtest):

typo\[\frac1{x^a}=x^{-a}\]

OpenStudy (anonymous):

so 1/x^1/2=x^-1/2

OpenStudy (turingtest):

right :) now what is\[\int x^ndx\]?

OpenStudy (anonymous):

would it be 2x^1/2

OpenStudy (anonymous):

+c